### v1

# Temperature Dependent Spin-Phonon Coupling of Boron-Vacancy Centers in Hexagonal Boron Nitride

Abstract
The negatively charged boron-vacancy center ($\mathrm{V}_{\mathrm{B}}^{-}$) in hexagonal boron nitride (hBN) has recently emerged as a highly promising quantum sensor. Compared to the nitrogen-vacancy (NV) center in diamond, the change with temperature of the spin transition energy of $\mathrm{V}_{\mathrm{B}}^{-}$ is more than an order of magnitude larger, making it a potential nanoscale thermometer with superior sensitivity. However, the underlying mechanism of the observed large temperature dependence remains an open question. In this work, using isotopically purified $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$, we systematically characterize the zero-field splitting, hyperfine interaction, and spin relaxation time of $\mathrm{V}_{\mathrm{B}}^{-}$ from 10 to 350 K. We carry out first-principle calculations of the $\mathrm{V}_{\mathrm{B}}^{-}$ spin-phonon interaction and show that a second-order effect from finite-temperature phonon excitations is responsible for the observed changes in experiments. By fitting our experimental results to a physically motivated model, we extract the dominant phonon mode which agrees well with our simulations. Finally, we investigate the dynamic nuclear spin polarization process at cryogenic temperatures. Our results provide key insights in $\mathrm{V}_{\mathrm{B}}^{-}$ centers and their utilization as nanoscale thermometers and phonon sensors.
Introduction— Optically-addressable solid state spin defects are promising platforms for quantum applications kolkowitz2012coherent ; doherty2013nitrogen ; aharonovich2016solid ; awschalom2018quantum ; atature2018material ; wolfowicz2021quantum ; togan2010quantum ; pompili2021realization ; degen2017quantum ; zu2021emergent ; koehl2011room ; nagy2019high ; hensen2015loophole ; randall2021many ; hsieh2019imaging ; thiel2019probing ; davis2023probing ; mittiga2018imaging ; he2023quasi ; dwyer2022probing ; Bhattacharyya2024-ko ; acosta2009diamonds ; de2021materials ; luo2024room , and when residing in atomically-thin van der Waals materials they may exhibit properties superior to those of their conterparts in three-dimensional materials azzam2021prospects ; ren2019review ; caldwell2019photonics ; naclerio2023review ; ye2019spin . Among a wide-range of contestants, the negatively charged boron-vacancy center, $\mathrm{V}_{\mathrm{B}}^{-}$, in hexagonal boron nitride (hBN) has attracted much growing interest gottscholl2020initialization ; gottscholl2021room ; gong2023coherent ; huang2022wide ; broadway2020imaging ; kumar2022magnetic ; healey2023quantum ; robertson2023detection ; gao2023quantum ; udvarhelyi2023planar ; hennessey2023framework ; patrickson2024high ; ivady2020ab ; kianinia2020generation . Several interesting properties of the host material, including the large bandgap and excellent stability, as well as the readily controllable spin degree of freedom at room temperature, makes $\mathrm{V}_{\mathrm{B}}^{-}$ in hBN well-suited as a platform for technologies.
When compared with its three-dimensional counterparts, for example the nitrogen-vacancy (NV) center in diamond, the ground state zero-field splitting (ZFS) of $\mathrm{V}_{\mathrm{B}}^{-}$ exhibits a $\sim 25$ times larger variation from cryogenic to room temperatures, placing it in an advantageous position for nanoscale thermometry vaidya2023quantum ; liu2021temperature ; gottscholl2021spin ; kucsko2013nanometre ; neumann2013high . However, the underlying mechanism of such a large temperature response remains an open question. In particular, the first-order lattice displacement of hBN with temperature can only account for a negligible fraction of the changes in the measured ZFS of $\mathrm{V}_{\mathrm{B}}^{-}$  gottscholl2021spin ; higher-order spin-phonon interactions should be included to quantitatively capture the experimental results.
In this letter, we experimentally characterize the temperature dependence of the ZFS, hyperfine interaction, and spin relaxation time $T_{1}$ of $\mathrm{V}_{\mathrm{B}}^{-}$ in isotopically purified $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$. Compared to conventional hBN with natural abundant 14N, our $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ offers much narrower $\mathrm{V}_{\mathrm{B}}^{-}$ spin transitions and substantially improved measurement resolution gong2024isotope ; janzen2023boron ; clua2023isotopic ; sasaki2023nitrogen . We then carry out first-principle calculations of the temperature-induced phonon-mediated interaction of $\mathrm{V}_{\mathrm{B}}^{-}$ , motivated by recent investigations of the NV centers in diamond cambria2023physically ; tang2023first . Our combined experimental and theoretical study allows for the identification of the main mechanism responsible for the observed temperature dependence of the properties of $\mathrm{V}_{\mathrm{B}}^{-}$, namely a second-order spin-phonon coupling with a characteristic phonon mode energy around 181818~{}18meV. We also investigate the dynamic nuclear spin polarization of the nearest three 15N and find that the polarization persists down to 101010~{}10K.
Figure 1: Temperature dependence of optically detected magnetic resonance spectra of $\mathrm{V}_{\mathrm{B}}^{-}$ (a) Schematic representation of a single $\mathrm{V}_{\mathrm{B}}^{-}$ center (red spin) in the hBN honeycomb lattice. $\hat{z}$ is defined as the out-of-plane direction, while $\hat{x}$ and $\hat{y}$ live in the lattice plane. (b) The $\mathrm{V}_{\mathrm{B}}^{-}$ electronic ground state energy level diagram with the presence of three nearest ${}^{15}\mathrm{N}$ nuclear spins in isotopically purified $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ flakes. The $|m_{s}=\pm 1\rangle$ is separated from $|m_{s}=0\rangle$ by a zero-field splitting $D$. The hyperfine interaction further divides each spin transition into four transitions with spacing $A_{zz}$ and degeneracy of {1, 3, 3, 1}. (c) ODMR spectrum of $\mathrm{V}_{\mathrm{B}}^{-}$ in $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ flakes under magnetic field $B_{z}\approx 90$ G at temperatures ranging from 10 K to 350 K. The normalized fluorescence (FL) contrast is marked in the colobar. The bottom panels of (c) and (d) display the ODMR spectrum at 300 K, corresponding to the black dashed line on the top panels. The spectrum is fitted with two groups of equally spaced Lorentzians to extract the values of ZFS and hyperfine splitting. (d) ODMR spectrum of $\mathrm{V}_{\mathrm{B}}^{-}$ in $\mathrm{hBN_{\mathrm{nat}}}$ flakes in the 10-350 K temperature range.
Experimental system— The $\mathrm{V}_{\mathrm{B}}^{-}$ is an atomic spin defect in the hBN lattice (Fig. 1a). The electronic ground state of $\mathrm{V}_{\mathrm{B}}^{-}$ exhibits a spin-1 triplet with $|m_{s}=0\rangle$ and $|m_{s}=\pm 1\rangle$. In the absence of external fields, the $|m_{s}=\pm 1\rangle$ are degenerated and separated from $|m_{s}=0\rangle$ by a zero-field splitting (ZFS) $D=(2\pi)\times 3.48~{}$GHz at room temperature (Fig. 1b).
The Hamiltonian governing $\mathrm{V}_{\mathrm{B}}^{-}$ and the nearest three nitrogen nuclear spins takes the form,
$\begin{split}\mathcal{H}&=DS_{z}^{2}+\gamma_{e}B_{z}S_{z}+\sum_{j=1}^{3}\mathbf{S}\mathbf{A}^{j}\mathbf{I}^{j}-\sum_{j=1}^{3}\gamma_{n}B_{z}I_{z}^{j},\end{split}$ (1)
where $B_{z}$ is an external magnetic field along the c-axis of hBN (or $\hat{z}$ direction), $\mathbf{S}$ and $S_{z}$ are the electron spin-1 operators, $\mathbf{I}^{j}$ and $I_{z}^{j}$ are the nuclear spin operators, and $\mathbf{A}^{j}$ is the hyperfine tensor coupling electronic and nuclear spins. The factors $\gamma_{e}=(2\pi)\times 2.8~{}$MHz/G and $\gamma_{n}$ are electronic and nuclear gyromagnetic ratios.
Throughout our experiment, the external magnetic field is $B_{z}\lesssim 760~{}$G. In this case, the energy splittings between $|m_{s}=0\rangle$ and $|m_{s}=\pm 1\rangle$ sub-levels, $D\pm\gamma_{e}B_{z}$, are much larger than the strength of the hyperfine interaction $|\mathbf{A}^{j}|$, thus the secular approximation can be adopted: $\sum_{j}\mathbf{S}\mathbf{A}^{j}\mathbf{I}^{j}\approx\sum_{j}A_{zz}^{j}S_{z}I_{z}^{j}=A_{zz}(\sum_{j}I_{z}^{j})S_{z}$, leading to a nuclear spin dependent energy shift of the electronic transition. This results in $12\mathcal{I}+$ 1 transitions between $|m_{s}=0\rangle$ and $|m_{s}=-1\rangle$ (or $|m_{s}=+1\rangle$) sub-levels, where $\mathcal{I}=\sum_{j}I^{j}$ is the total nuclear spin number accounting for all three nearest nitrogen atoms (Fig. 1b).
The corresponding spin transitions of $\mathrm{V}_{\mathrm{B}}^{-}$ can be probed via optically detected magnetic resonance (ODMR) spectroscopy: by sweeping the frequency of the applied microwave drive while monitoring the fluorescence signal of $\mathrm{V}_{\mathrm{B}}^{-}$, we expect a fluorescence drop whenever the microwave is resonant with an electronic spin transition. We report measurements for hBN with naturally distributed isotopes $\mathrm{hBN_{\mathrm{nat}}}$ (99.6% ${}^{14}\mathrm{N}$ with $I=1$), and for the isotope-engineered $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ (99.7% ${}^{15}\mathrm{N}$ with $I=\frac{1}{2}$gong2024isotope ; janzen2023boron . For conventional $\mathrm{hBN_{\mathrm{nat}}}$, there are in total seven broad hyperfine spin transitions; while for $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$, we observe four sharp resonances, enabling the measurement of ZFS and hyperfine splitting, $A_{zz}$, with substantially improved resolution (Fig. 1cd).
Experiment results— To investigate the temperature-dependent spin properties of $\mathrm{V}_{\mathrm{B}}^{-}$, we load hBN samples with thickness ranging from $\sim 30-70$ nm into a closed-cycle optical cryostat (Fournine Design) for temperature control from $10-350~{}$K. A small external magnetic field $B_{z}\approx 90$ G is used to lift the degeneracy between $|m_{s}=\pm 1\rangle$. Figure 1c,d show the ODMR spectra as a function of temperature. We fit the spectra with two groups of equally spaced Lorentzians, corresponding to $|m_{s}=\rangle\leftrightarrow|m_{s}=\pm 1\rangle$ transition respectively. The ZFS, $D(T)$, can be extracted from the average frequency of the two transition groups, while the hyperfine splitting, $A_{zz}(T)$, is obtained from the fitted spacing between adjacent resonances in each group (Fig. 2a). In addition, we also perform spin relaxation measurements and extract the $T_{1}$ timescale of the sample under different temperatures (Fig. 2b).
Figure 2: Temperature-dependent properties of $\mathrm{V}_{\mathrm{B}}^{-}$ in different hBN samples. (a) The temperature dependence of the ZFS $D(T)$ of $\mathrm{V}_{\mathrm{B}}^{-}$ in the range 10-350 K. The solid lines represent a fit to a physically motivated model using Eq. 3. Inset: hyperfine interaction $A_{zz}(T)$ of $\mathrm{V}_{\mathrm{B}}^{-}$ in $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ within the same temperature range, fitted by the same model with one fixed phonon energy $18.4$ meV extracted from the fit of the ZFS. (b) Spin relaxation rate $1/T_{1}$ of $\mathrm{V}_{\mathrm{B}}^{-}$ in the temperature range 10-350 K. Setting the phonon energy at $\hbar\omega_{\mathrm{exp}}$, the dotted lines qualitatively reproduce the $T_{1}$ temperature-dependence by the model Eq. 4. In the high temperature regime, the relaxation rate is approximated by a power-law scaling $1/T_{1}\propto T^{2}$. The experimental $T_{1}$ pulse sequence is given in the inset.
We find that the ZFS of $\mathrm{V}_{\mathrm{B}}^{-}$ exhibits a dramatic change of more than $200$ MHz from 10 to 350 K, which is around 30 times larger than the change observed for NV centers cambria2023physically . We determine the susceptibility of $\mathrm{V}_{\mathrm{B}}^{-}$ at room temperature, $\chi(T=300~{}\mathrm{K})$, to be $-880\pm 65$ kHz/K for $\mathrm{hBN_{\mathrm{nat}}}$, and $-784\pm 35$ kHz/K for $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$. These results highlight the potential use of $\mathrm{V}_{\mathrm{B}}^{-}$ as an ultra-sensitive nanoscale thermometer. Furthermore, within our measured temperature range, we observe a slightly smaller change of ZFS in $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ compared to $\mathrm{hBN_{\mathrm{nat}}}$, which can be ascribed to the fact that the heavier nuclei in $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ lead to atomic displacements less sensitive to the temperature change. The isotope effect is further evidenced by the measured $T_{1}$ timescales, limited by spin-phonon interaction gottscholl2021room ; durand2023optically ; cambria2023temperature : for heavier nuclei and thus weaker spin-phonon coupling strength, the $T_{1}$ of $\mathrm{V}_{\mathrm{B}}^{-}$ in $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ is further extended at lower temperatures. We also note that while it remains a challenge to resolve the hyperfine interaction strength from the broad ODMR spectra of conventional $\mathrm{hBN_{\mathrm{nat}}}$  gong2024isotope , the substantially better resolved resonances of $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ enable an accurate characterization of $A_{zz}(T)$ (see Supplementary Materials). The measured amplitude of $A_{zz}(T)$ displays a noticeable increase from 646464~{}64MHz to 676767~{}67MHz with decreasing temperature, whereas the hyperfine interaction of NV centers to the nearby nuclear spins has been measured to be nearly constant with temperature barson2019temperature ; xu2023high . Finally, we observe that the ODMR contrast of $\mathrm{V}_{\mathrm{B}}^{-}$ in both hBN samples reaches a maximum around 210210210~{}210K and persists to low temperatures, without any substantial quenching. This behaviour is again in opposite to that of the NV center in diamond, whose contrast suffers from a significant drop below 909090~{}90fischer2013optical ; ernst2023temperature ; happacher2023temperature , hindering its sensing performance at cryogenic temperatures.
Table 1: Zero-field splitting fitting parameters. For each isotope, the extracted parameters are averaged over four sets of data in different sample spacial positions (see Supplementary Materials), fitted with same model in Eq. 3. We note that $D_{0}$ differs from $D(T=0)$, where $D_{0}=D(T=0)-\frac{1}{2}c_{D}$.
Isotopes $D_{0}$ (MHz) $c_{D}$ (MHz) $\hbar\omega_{\mathrm{exp}}$ (meV)
${}^{15}\mathrm{N}$ $3742\pm 10$ $-175\pm 13$ $18.4\pm 1.0$
${}^{14}\mathrm{N}$ $3777\pm 26$ $-201\pm 42$ $18.8\pm 2.8$
Theoretical model— To understand the observed temperature dependence of $\mathrm{V}_{\mathrm{B}}^{-}$ spin properties, we adopt a theoretical model which was originally developed to study NV centers in diamond tang2023first . In particular, we write the electronic and nuclear spin transitions as the sum of two terms,
$\nu=\nu_{0}(a(T))+\frac{1}{2}\sum_{i}\frac{\partial^{2}\nu}{\partial q_{i}^{2}}\frac{\hbar}{M_{i}\omega_{i}}\left(\frac{1}{e^{\hbar\omega_{i}/k_{B}T}-1}+\frac{1}{2}\right),$ (2)
where $\nu$ is the transition frequency (e.g., ZFS and hyperfine interaction) and $a(T)$ is the lattice constant of hBN as a function of temperature gottscholl2021spin , $q_{i},M_{i},\omega_{i}$ are the normal mode of the crystal, mode-specific effective mass, and frequency, respectively. Here the first-order term, $\nu_{0}(a(T))$, corresponds to the thermal expansion of the hBN lattice, while the second-order term represents vibrational contributions caused by finite-temperature phonon excitations near the equilibrium geometry.
We carry out first-principle simulations using density functional theory (DFT) kohn1965self with the Perdew-Burke-Ernzerhof (PBE) functional perdew1996generalized to compute the temperature-dependent hyperfine interaction, $A_{zz}(T)$, between $\mathrm{V}_{\mathrm{B}}^{-}$ and the nearest three ${}^{15}\mathrm{N}$ nuclei. Interestingly, we find that the variation of the lattice constant with temperature only results in an extremely small change in the magnitude of $A_{zz}(T)$. On the other hand, the second-order term accounts for the effect of temperature on the value of $A_{zz}(T)$ observed experimentally.
[…]
Figure 3: (a) The second-order vibrational coupling, i.e., $\frac{\partial^{2}A_{\mathrm{zz}}}{\partial q_{i}^{2}}\frac{\hbar}{M_{i}\omega_{i}}$ as a function of phonon frequencies of the nearest ${}^{15}\mathrm{N}$ in hBN (blue) and 13C in diamond (orange) to the vacancy, respectively. The first peak of $\mathrm{V}_{\mathrm{B}}^{-}$ in hBN is identified as an out-of-plane vibrational mode (inset). (b) The computed $A_{\mathrm{zz}}(T)$ of the nearest ${}^{15}\mathrm{N}$ in hBN (blue) from both the first and second order vibrational contribution, respectively. We also plotted the computed $A_{\mathrm{zz}}(T)$ of the nearest 13C in NV- in diamond (orange) from the second-order vibrational contribution as a comparison.
The dominant phonon mode found theoretically at $\hbar\omega_{\mathrm{th}}\approx 16~{}$meV can also be independently determined by fitting the experimentally measured spin properties to a physically motivated model cambria2023physically ,
$\nu(T)=\nu_{0}+c_{\nu}\left(\frac{1}{e^{\hbar\omega/kT}-1}+\frac{1}{2}\right),$ (3)
where $\nu(T=0)=\nu_{0}+\frac{1}{2}c_{\nu}$ is the transition energy (e.g. ZFS $D$ and hyperfine $A_{zz}$) at 0 K and $n=(e^{\hbar\omega/kT}-1)^{-1}$ is the occupation number of the phonon mode.
We apply this model to the measured ZFS and hyperfine interaction of $\mathrm{V}_{\mathrm{B}}^{-}$. For ZFS, the observed temperature dependence of $D(T)$ for both $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$ and $\mathrm{hBN_{\mathrm{nat}}}$ samples can be accurately fitted using our model, from which we extract the characteristic phonon energy to be $\hbar\omega_{\mathrm{exp}}=18.4\pm 1.0~{}$meV and $18.8\pm 2.8~{}$meV respectively (Fig. 2a and Table 1), in agreement with our theoretical results. The measured hyperfine interaction, $A_{zz}(T)$, shows a temperature dependence that is also consistent with the model with phonon energy set to $18.4$ meV (Fig. 2a Inset). Interestingly, the agreement between $\hbar\omega_{\mathrm{exp}}$ and the calculated $\hbar\omega_{\mathrm{th}}$ suggests that the temperature dependence of ZFS and of the hyperfine interaction originate from the coupling of the electronic spin to the same vibrational phonon mode.
We now turn to investigate the temperature dependence of the spin relaxation time $T_{1}$. Within the temperature range ($10-350$ K) considered here, $T_{1}$ of $\mathrm{V}_{\mathrm{B}}^{-}$ is dominated by Raman scattering processes, where the energy difference between phonon absorption/emission is equal to the ZFS of $\mathrm{V}_{\mathrm{B}}^{-}$, leading to spin depolarization (see Supplementary Materials) cambria2023temperature . Therefore the spin relaxation rate of $\mathrm{V}_{\mathrm{B}}^{-}$ can be approximated as cambria2023temperature
$\Gamma(T)=1/T_{1}\approx An(n+1)+A_{S},$ (4)
where $n=(e^{\hbar\omega/kT}-1)^{-1}$ and $A$ are the phonon occupation number and coupling coefficient associated with the effective mode, and $A_{S}$ is a sample-related constant that might vary slightly between experiments. Using the phonon energy $\hbar\omega_{\mathrm{exp}}=18.4\;(18.8)~{}$meV determined above, the model of Eq. 4 faithfully reproduces the experimentally measured spin relaxation rates from $50-350~{}$K (Fig. 2b). The small discrepancy below 505050~{}50K may be due to the activation of lower-energy phonon modes as well as processes beyond Raman scattering. We note that a prior theoretical study has predicted a power-law scaling of $T_{1}\sim T^{-2}$ mondal2023spin , which is consistent with our experimental results in the high temperature regime (Fig. 2b).
Dynamic nuclear polarization at low temperature— With the temperature-dependent $\mathrm{V}_{\mathrm{B}}^{-}$ spin properties in hand, we now turn to investigate the dynamical polarization of the three nearest-neighbor ${}^{15}\mathrm{N}$ nuclear spins at low temperature. Nuclear spins feature exceptional isolation from noisy environments, making them ideal candidates for quantum storage applications bradley2019ten ; metsch2019initialization . At room temperature, several prior works have demonstrated the polarization of proximate 14N or 15N with an upper limit of $\sim 30~{}\%$; however how to further improve the polarization fidelity remains an open question  gong2024isotope ; gao2022nuclear ; ru2023robust . Low temperature offers two potential advantages: $\sim 100$x longer $\mathrm{V}_{\mathrm{B}}^{-}$ electronic spin lifetime $T_{1}$ and stronger hyperfine interaction strength.
To enable resonant spin exchange between electronic and nuclear spins (known as electronic spin level anti-crossing, esLAC) gong2024isotope ; gao2022nuclear ; mathur2022excited , we apply an external magnetic field $B_{z}\approx 760~{}$G under which the energy levels corresponding to $\mathrm{V}_{\mathrm{B}}^{-}$ excited state $|m_{s}=0\rangle$ and $|m_{s}=-1\rangle$ are nearly degenerate (see Supplementary Materials). We fit the measured hyperfine resonances between $\mathrm{V}_{\mathrm{B}}^{-}$ $|m_{s}=0\rangle\leftrightarrow|m_{s}=+1\rangle$ in the ODMR spectra as a function of the laser power to extract the nuclear spin polarization fidelity (Fig. 4a). We find that the nuclear spin polarization at both $10$ K and 300300300~{}300K saturate around $30\%$ with similar laser powers ($\sim 10~{}$mW) (Fig. 4b), suggesting that the $\mathrm{V}_{\mathrm{B}}^{-}$ spin lifetime $T_{1}$ and hyperfine interaction strength are not the limiting sources for nuclear polarization.
Figure 4: Dynamic nuclear polarization at low temperature (a) Near esLAC level ODMR Spectra of the $|m_{s}=0\rangle\leftrightarrow|m_{s}=-1\rangle$ transition at 10 K and 300 K. The spectra both exhibit similar asymmetry toward the left peaks, indicating a polarization of nuclear spins (see Supplementary Materials) (b) Polarization level of the three nearest ${}^{15}\mathrm{N}$ nuclear spins near esLAC level as a function of the normalized laser power under room/cryogenic temperature. Inset: schematic representation of the polarization process. The strong spin-conserving optical polarization (green arrow) continuously pumps state from $|m_{s}=-1\rangle$ to $|m_{s}=0\rangle$, and the two hybridisation processes (red arrow and dashed blue arrow) differ in strength, resulting in the polarization of $|m_{I}=\>\uparrow\rangle$ state (see Supplementary Materials).
The spin polarization may be limited from the spin non-conserving terms in the hyperfine interaction (see Supplementary Materials), which are temperature insensitive. In addition, the observed nuclear spin polarization of $\mathrm{V}_{\mathrm{B}}^{-}$ persists down to 101010~{}10K, in sharp contrast to the case of NV centers where the dynamic nuclear polarization process is substantially suppressed at $T\lesssim 50~{}$K, due to the lack of thermal averaging between NV excited states originating from different orbitals fischer2013optical ; block2021optically . For $\mathrm{V}_{\mathrm{B}}^{-}$ , the lifting of the double degeneracy of excited states leads to a sufficiently large energy gap between the states, which renders the system insensitive to dynamic Jahn–Teller distortions mathur2022excited .
Conclusion and Outlook— In summary, our research offers a comprehensive investigation of the temperature-dependent spin-phonon interaction of $\mathrm{V}_{\mathrm{B}}^{-}$ centers in isotopically purified $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$. Our results will be of critical importance in designing experiments to make use of $\mathrm{V}_{\mathrm{B}}^{-}$ centers as nanoscale thermometers and phonon sensors. Moreover, although we focus on $\mathrm{V}_{\mathrm{B}}^{-}$, the methodologies developed here can be readily applied to a broad family of spin defects in two-dimensional materials stern2023quantum ; guo2023coherent ; li2022carbon ; lee2022spin .
Acknowledgements: We gratefully acknowledge Vincent Jacques, Tongcang Li, Shankar Mukherji, Erdong Song, Xi Wang and Norman Yao for helpful discussions. This work is supported by NSF ExpandQISE 2328837 and the Center for Quantum Leaps of Washington University. B. Huang, Y. Jin and G. Galli acknowledge the support of Air Force Office of Scientific Research (AFOSR) under Contract No. FA9550-22-1-0370, the Next Generation Quantum Science and Engineering (QNEXT) hub supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers and resources of the University of Chicago Research Computing Center. Support for hBN crystal growth (E. Janzen and J. H. Edgar) was provided by the Office of Naval Research, award no. N00014-22-1-2582.